OpenStudy (anonymous):

5i(13 - 8i)...I keep getting 65 - 40i

OpenStudy (espex):

You mean you get 65i+40?

OpenStudy (anonymous):

no, I keep getting 65 - 40i, and I know its wrong but I don't know what I am doing wrong?

OpenStudy (espex):

5i(13-8i) => 65i-40i^2 => i^2 = -1 => 65i-(-1)(40) => 65i+40

OpenStudy (espex):

Is that not the way you are doing it?

OpenStudy (anonymous):

nope. I'm just multiplying 5i by 13, then by 8i

OpenStudy (espex):

Exactly, so 5i * 13 = 65i 5i * 8i = 40(i^2) i^2 = -1

OpenStudy (anonymous):

oh I see now

OpenStudy (espex):

Glad to hear it.. :)